______     2.      Assume a firm would like to borrow $900 from a bank for one year to invest in a project with the following possible cash flows at time 1:

Probability

40% chance

60% chance

Project cash flows

$375

$2000

The firm has no “assets-in-place,” and therefore will only be able to use the cash flows from the project (given above) to make payment on the loan. What interest rate will the bank need to charge on this one-year loan to earn an expected return of 7%?

A. 36.28%

B. 39.44%

C. 41.10%

D. 43.15%

E. 46.85%

F. 48.23%

G. 50.56%

H. 52.17%
Answer #1

900*1.07=40%*375+60%*x
=>x=(900*1.07-40%*375)/60%
=>x=$1,355.00

Interest rate the bank need to charge=1355/900-1=50.556%